BG February Com Night 4, February 4 2009

Observers: Andrew Rakich
Telescope Operator: David Gonzelez
SW Support: Name
Telescope Support: Name

Summary:

Got some good field measurements. Seeing was ~ 1" throughout, but active optics was converging much better. Had a few issues with AGW and GCS see in the log below.

Preparations:

Continue with the program started last night. (More field aberration measurements at the end of MCS work).

Open

Details (Times in UT):

10:30 setting up to do field measurements, collimating on BS9145 GS 0. Adjust IE=-44 CA=10

10:32 wind 3.6 m/s Temp 1.7 deg C seeing ~ 1"-1.5"

*BS9145 - Field Aberration Measurements
GStar X Y PA Doug's AngleSorted ascending wfscimage000
0 0.176 612.391 343.83 -534 26-36
6 -0.858 477.486 343.83 -531 20-22
0 -0.132 612.759 194.32 -394 47- 51
7 -1.011 479.094 194.32 -393 44-46
0 -0.103 612.74 194.32 -379 16-19
7 -1.068 479.07 194.32 -377 12-14
0 0.025 612.418 120.92 -327 61-68
19 -0.560 467.516 120.92 -325 58-60
0 -0.279 612.519 120.92 -318 40-43
19 -0.538 467.536 120.92 -317 37-39
0 0.147 612.397 343.83 -187 55-57
6 -0.813 477.496 343.83 -186 52-54
0 0.044 612.352 0.00 -180 2-11

11:45 noticed some Z22 on the primary ~ 100 nm and the temp and dropped 1 deg C in 1 hour. drove it out manually in psf (-300 nm Z22 in psf made the measured SH Z22 drop by 100 nm).

11:53 go to BS9152, tried to send ACQUIRE preset over the top of guide preset and got "error Newline while processing string" Killed GCS, didn't help, killed IRAF and restarted. adjust pointing IE= -40 CA=6

*BS9152 - Field Aberration Measurements
GStar X Y PA Doug's Angle wfscimage000
0 -0.147 612.623 0.00 -180 70-72
9 -1.221 480.415 168.12 -350 73-75
0 -0.147 612.424 168.12 -352 76-77
17 -0.123 436.83 124.77 -311 78-80
0 -.073 612.364 124.77 -312 81-83
18 -1.292 426.590 65.43 -255 84-86
0 .073 612.429 65.43 -256 87-89
7 -0.529 501.177 10.54 -206 90-92
0 .117 612.510 10.54 -207 93-94
4 -0.913 476.024 338.61 -179 95-97
0 .117 612.680 338.61 -180 98- 100
1 -0.901 501.381 229.72 -433 101-103
0 -0.162 612.624 229.72 -434 104-107
18 -1.298 426.540 65.43 -282 108-110
0 0.103 612.379 65.43 -283 111-112

12:45 Torsten, several times, this time (5:45 MST) was the latest, when a star was borderline too faint, the acquisition would fail, then resending the preset to the same star would succeed.

1:06 (6:06 MST) Torsten, had a star in the guide box, then guiding threw it right out of the box and preset failed, then GCS hung. Seems that the elevation axis got stuck which was our problem. GCS failed to report this source of error (some error message like PCS didn't return or whatever would be useful).

1:33 AGW failed to deliver an image to GCS. Note the UT clock in the control room has lost ~5 minutes since last night when I reset it (it was 30 minutes behind then). New battery?

Close

HH:MM Close dome T=?degC, D=?degC, wind (LBTO) ? m/sec.

-- JoarBrynnel - 15 Jan 2009
Topic revision: r6 - 05 Feb 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback